Grade 11Physical Chemistry

X gm A atoms on combining with Y atoms of B form 5 molecules of a compound containing A and B. find the molecular weight of compound formed. (atomic weight of B =M)

Kashish Prajapati
8 Years agoGrade 11

1 Answer

ADARSH SINGH
8 Years ago
AS THIS INFORMATION IS NOT GIVEN HERE , WE ARE CONSIDERING THAT ALL AMOUNT OF SUBSTANCE A REACTS WITH GIVEN ATOMS OF B TO FORM COMPOUND CONTAINING A AND B.
NOW, TOTAL MASS OF A REACTED= Xgm
TOTAL MASS OF B REACTED = Y x M gm (no. of atoms x atomic weight)
total mass of reactans = X + YM gm
BY LAW OF CONSERVATION OF MASS, TOTAL MASS OF REACTANT IS EQUAL TO TOTAL MASS OF THE PRODUCTS .
THERE FORE MASS OF 5 MOLECULES OF FORMED COMPOUND = X + YM
THEREFORE MASS OF 1 MOLECULE OF GIVEN COMPOUND FORMED = (X + YM)/5